8a^3+c^3=
i really dont knowhow to do this

Answer by lwsshak3(11628)

8a^3+c^3=2^3a^3+c^3
=(2a)^3+c^3
use following sum of cubes formula to solve:
x^3+y^3=(x+y)(x^2-xy+y^2)
For given problem:
x=2a
y=c
You can take it from here to solve.